OpenStudy (anonymous):

Factor the polynomial completely. (If it is not factorable using integers, enter NF.) x^4 -13 x^2+36

OpenStudy (anonymous):

\[(x-3) (x-2) (x+2) (x+3) \]

OpenStudy (anonymous):

\[x^{4}-13x^{2}+36=\left( x^{2}-9 \right)\left( x^{2}-4 \right)=\left( x+3 \right)\left( x-3 \right)\left( x+2 \right)\left( x-2 \right)\]